What is the amount of 10 equal annual deposits that can provide five annual withdrawals
when the first withdrawals of $5,000 is made at the end of Year 11 and each subsequent
withdrawal increases by $400 if the interest is 9 percent compounded quarterly?
1
Expert's answer
2020-10-23T07:21:32-0400

1.PV of all withdrawals:

11 5000 1878,3827

12 5400 1855,8999

13 5800 1823,625

14 6200 1783,3885

15 6600 1736,7805

9078,0765


"\frac{5000}{(1+\frac{0.09}{4})^{11*4}}=1878.38"


And so on, add 400 and find PV

sum is 9078.08

2.

we will calculate the divide for all deposit periods:

cells F to T





  1. find 10 equal annual deposits 

"\frac{x\times(1-0.410646)}{(1-0.9148)}=9078.08"

x=1311.71